OpenStudy (loser66):

How to prove \(X_n = \dfrac{n^2}{n+1}\) diverge?

OpenStudy (mathmath333):

using limit ?

OpenStudy (loser66):

Using limit theorem, not definition.

OpenStudy (loser66):

Limit theorems talk about convergent sequence only. That is the problem, :(

OpenStudy (anonymous):

Use the fact that \(\dfrac{n^2}{n+1}=n-1+\dfrac{1}{n}\).

OpenStudy (loser66):

Yes, I did, but lim n = infinitive get 0 credit from my prof. :)

OpenStudy (loser66):

He said: you can break X_n into many parts as you like, but each part must be a convergent to apply limit theorem there. Your n is a divergent sequence and we never construct the formula lim n = infinitive.

ganeshie8 (ganeshie8):

try assuming there exists a limit L and arrive at a contradiction

ganeshie8 (ganeshie8):

\[\large \left|\frac{n^2}{n+1} - L\right|\lt \epsilon\]

ganeshie8 (ganeshie8):

or are you not allowed to use contradiction ?

ganeshie8 (ganeshie8):

lol he gave you negative marks :o

OpenStudy (loser66):

Yes. :)

OpenStudy (anonymous):

can't you use L'Hopital's rule

OpenStudy (loser66):

:( No, I can't. Just limit theorem. even definition of limit is not allowed. :(

OpenStudy (loser66):

this page is next to the previous one. Just take a look for fun. :)

OpenStudy (anonymous):

:S

ganeshie8 (ganeshie8):

\[\large \lim\limits_{n\to \infty} \dfrac{n^2}{n+1} =\lim\limits_{n\to \infty} \dfrac{n^2}{n} = \lim\limits_{n\to\infty} n =\infty \]

ganeshie8 (ganeshie8):

because as n->infty, n+1 ~ n

OpenStudy (loser66):

Ok, hopefully with this logic, I can get back 2 credits. hahaha.. No worry, I am already get A^- for this homework. :)

ganeshie8 (ganeshie8):

idk if there is a more pleasing proof @nerdguy2535

OpenStudy (loser66):

he is offline :)

ganeshie8 (ganeshie8):

@freckles @zepdrix @AccessDenied

OpenStudy (freckles):

You should able to prove just lim n-> inf n=inf since \[\frac{n^2}{n+1}=n+\frac{-1}{n+1}\] and -1/(n+1) goes to 0

OpenStudy (anonymous):

I think the idea that Sith had above can work, you just have to work it properly. We have that: $$\frac{n^2}{n+1}=n-1+\frac{1}{n}\iff n=\frac{n^2}{n+1}-\frac{1}{n}+1.$$We proceed by contradiction. Assume that: $$\lim_{n\rightarrow \infty}\frac{n^2}{n+1}=L.$$ Since: $$\lim_{n\rightarrow \infty }\frac{1}{n}=0$$and: $$\lim_{n\rightarrow \infty }1=1,$$using the fact that the limit of the the sum of three convergent series is the sum of the limits, we see that: $$\lim_{n\rightarrow \infty }n=\lim_{n\rightarrow \infty }\left(\frac{n^2}{n+1}-\frac{1}{n}+1\right)=\lim_{n\rightarrow \infty }\left(\frac{n^2}{n+1}\right)-\lim_{n\rightarrow \infty }\left(\frac{1}{n}\right)+\lim_{n\rightarrow \infty }\left(1\right)$$$$=L-0+1=L+1,$$which is absurd since the sequence \(\{n\}_{n=1}^\infty\) is unbounded.

OpenStudy (loser66):

Thank you so much.